OpenStudy (laughoutloud):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3> -9.1° 1.8° 0.9° 11.8°

OpenStudy (jabberwock):

The first thing to know is the dot product between two vectors. Are you familiar with the dot product?

OpenStudy (jabberwock):

\[v \cdot u=||v|| ||u||\cos(\theta)\] You can calculate the dot product of the vectors (on the left side). You can calculate the magnitudes of the vectors on the right side (using the Pythagorean Theorem). Isolate the cos(theta) Then you can take the inverse cosine to solve for theta.

OpenStudy (mathmale):

Yes, this is precisely the correct approach. Laugh: Please solve this equation for cos theta. type out your result here.

OpenStudy (mathmale):

Further questions welcome, if you need clarification.